A proton is released from rest at the positive plate of a parallel plate capacitor. The charge per unit area on each plate is σ=1.8e-7 C/m2 , and the plates are separated by a distance of 1.5e-2 m.

a. What is the speed of the proton when it reaches the negative plate? Solve this problem using conservation of energy.

b.Solve (a) using kinematics equations.
Answer #1

Electric field between the plates of the capacitor E=rac{sigma}{epsilon_0} directed from positive plate to negative plate.

Potential difference between the plates of the capacitor is V=Ed=rac{sigma d}{epsilon_0} ( if the +ve plate is potential V then -ve plate is at potential 0)

a)

Initial kinetic energy of proton 2 2

Initial potential energy of proton σ de

Final kinetic energy of proton K_f=rac{1}{2}mv^2

Final potential energy of proton P_f=0

Conserving the energy of proton,

K_i+P_i=K_f+P_f

ơde 0

2ơde Eom

2 18 10-1,510 1.6 10-19 8.85 1011.67 102 = 2.42 * 10° m/s V)

------

b) Force on the proton in the electric field E is  F=eE

ma=eERightarrow a=rac{eE}{m}

Using kinematics equation  v^2-u^2=2ad Rightarrow a=rac{v^2-u^2}{2d}

rac{v^2-u^2}{2d}=rac{eE}{m}

v=sqrt{rac{2deE}{m}+u^2}=sqrt{rac{2desigma}{epsilon_0m}+u^2}

v=sqrt{rac{2*1.5*10^{-2}*1.6*10^{-19}*1.8*10^{-7}}{8.85*10^{-12}*1.67*10^{-27}}+0^2}=2.42*10^5,m/s

------------------------------